OpenStudy (anonymous):

What is the surface area of a die in which each edge has a length of 15 millimeters? (Hint: On a cube, the length, width, and height have the same measure.) 1,350 mm2 450 mm2 180 mm2 1,687.5 mm2

jimthompson5910 (jim_thompson5910):

What is the area of one face of the cube?

OpenStudy (anonymous):

15mm

jimthompson5910 (jim_thompson5910):

that's one side, not the area

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

Area of one face = ???

OpenStudy (anonymous):

So you would times 15 by 2?

jimthompson5910 (jim_thompson5910):

Area = base * height A = b*h A = 15*15 A = 225

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

There are 6 of these faces in a cube (look at any dice) so total surface area is 225*6 = 1350 sq mm

OpenStudy (anonymous):

Thank you ! I see now!

jimthompson5910 (jim_thompson5910):

you're welcome